# Safe MPC Alignment with Human Directional Feedback

### v2

@@ IV-C Properties and Geometric Interpretations
In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment. Those properties will be used to establish the convergence of the algorithm in the next section.
IV-C1 Piecewise Linearity of Cuts
By recalling the weight-feature parameterization of the safety constraint (5), one can examine that the inequities in (12) are both linear in $\theta$. This means that the cutting set $\mathcal{C}_{i}$ in (12) is a piecewise linear cut, as stated formally in the following lemma.
@@ IV-C1 Piecewise Linearity of Cuts
[…]
@@
According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which is two hyperplanes:
@@
shown in Fig. 4.
As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplane pairs yields a series of convex polytopes in $\mathbb{R}^{r}$. Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
@@ IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
According to (14), the cutting set $\mathcal{C}_{i}$ depends on plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ which is a solution to penalty-based MPC (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The following assertion states a relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$.
@@
We briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig. 5 and Fig. 5, the removal of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$. Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal volume of the hypothesis space can be relatively small, as in Fig. 5, or large, as in Fig. 5, depending on the specific direction of the correction.
As the specific direction of user correction is hard to assume, one can intuitively see that a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$. As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig. 5. We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee of volume reduction rate of the hypothesis space in the next section.
Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$. (a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removal volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$. (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
V Main Results and Implementation Details
In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each learning iteration $i$ in our Safe MPC Alignment algorithm. Based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate in the hypothesis space volume for all possible human directional corrections. The reduction rate enables us to establish an exponential convergence of the learning process. As a result, we can find an upper bound of total human corrections for successful learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
@@ V-A The Choice of $\boldsymbol{\theta}_{i}$
As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in each learning iteration. One option is to choose the mass center of $\Theta_{i-1}$. However, computing the centroid of a polytope can be #P-Hard [53]. While methods like the Hit-And-Run algorithm [54] estimate the mass center by sampling, they can suffer from high sample complexity. Another possible choice is to use the center of the maximum inscribed ball of a polytope [55], but it lacks of the guarantee in volume reduction rate. For computational efficiency and guaranteed convergence, we will choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$, shown in Fig. 6.
Figure 6: $\boldsymbol{\theta}_{i}$ is chosen as the center of MVE of $\Theta_{i-1}$ (2D example).
Definition 1 (Maximum Volume inscribed Ellipsoid (MVE)[56]).
@@
=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
@@
where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$. Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [56]:
@@
Recall that in our Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by a pair of hyperplanes (17), and that the initial hypothesis space (11) is a box. By induction, it leads to that any updated hypothesis space $\Theta_{i}$ is a convex polytope, $i=0,1,2,...$. Thus, for any $\Theta_{i-1}$, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$. Then, the choice of $\boldsymbol{\theta}_{i}$ is
@@
The convex optimization (20) can be solved efficiently [57].
@@ V-B Result 1: Finite Iteration Convergence
In the Safe MPC Alignment algorithm, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
@@
The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume. With this, we can now establish the finite iteration convergence of the Safe MPC Alignment, stated below.
@@
The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$. In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety constraint intended by the user. The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$. Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$. It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated of $\boldsymbol{\theta}_{H}$ learned at convergence.
However, the finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm. Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a guaranteed method to detect the hypothesis misspecification, where such condition is not met. Next, we will discuss the declaration of hypothesis misspecification.
V-C Result 2: Hypothesis Misspecification Declaration
We consider the case of hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$. In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4). As shown in [52, 58], the hypothesis misspecification can lead to the undesired behavior of the robot. The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
@@ V-C Result 2: Hypothesis Misspecification Declaration
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
@@
The above result states that when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
The above theorem inspires us to empirically identify the misspecification by monitoring the position of the MVE center of $\Theta_{i}$ during the learning update. As shown in Fig. 7, if after the maximum iteration $K-1$ in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$. So we can use the distance of $\boldsymbol{\theta}_{K-1}$ to the boundary as an indicator for misspecification.
Specifically, recall that $\Theta_{0}$ in (10) is a box with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$. If $\boldsymbol{\theta}_{K-1}$ satisfies
@@ V-D Detailed Algorithm of Safe MPC Alignment
With the main results on convergence and misspecification, we can establish the full safe MPC alignment algorithm in Algorithm 1. With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24). Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3. After $K-1$ iterations, the algorithm checks the misspecification by (25).
Initialization: Parameterized constraint $\boldsymbol{g}_{\theta}$, initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$ (i.e., approximate radius of $\bar{\Theta}_{H}$), $i=1$
Compute maximum correction count $K$ from $\rho_{H}$ in (24);
@@
             Human restarts robot from safe region.;
@@
VI Numerical Examples and Analysis
In this section, we will evaluate the proposed Safe MPC Alignment using two numerical examples: inverted pendulum and two-link torque-controlled robot arm moving in constrained environment. Simulated human corrections knowing ground-truth safety constraints will be used. We will compare the proposed method with two baselines, and conduct ablation studies on the hyperparameters of the algorithm. All numerical experiments are implemented in Python, running on a Dell desktop with Intel i9-13900K CPU. All our experiment codes are open-sourced at at https://github.com/asu-iris/Safe-MPC-Alignment.
@@ VI-A1 Setup
with $\alpha$ being the angle between the pendulum and gravity, $\mathrm{g}$ the gravity constant, $u$ input torque, $m$, $l$ and $d$ the length, inertia, and damping ratio. We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (26) by the Euler method with $\triangle t=0.02\mathrm{s}$. A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each control step to simulate the state noise. The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$. We additionally pose an unknown global state constraint on pendulum motion:
@@
respectively, with $R=0.1$ and $Q=\mathrm{diag}(25,10)$. To ensure the safe MPC policy to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state. Thus, in (1), we define
@@
Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$. The simulated human corrections $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first entry is taken considering (6)). The correction is probabilistic to simulate the randomness of human providing corrections. It is easy to verify that the simulated $\mathbf{a}$ is valid in term of satisfying the hypotheses in section IV-A.

### v3

@@ IV-C Properties and Geometric Interpretations
We next show properties and geometric interpretations for the proposed safe MPC alignment process. Those properties will be used to establish the convergence of the algorithm.
IV-C1 Piecewise Linear Cut
By recalling the weight-feature parameterization of the safety constraint (5), one can examine that the inequities in (12) are both linear in $\theta$. This means that the cutting set $\mathcal{C}_{i}$ is a piecewise linear cut, as stated below.
@@ IV-C1 Piecewise Linear Cut
$b_{i}=\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})-\gamma\left<\mathbf{\bar{a}}_{i},\nabla\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right$>. (16)
@@
According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with two hyperplanes (Fig. 4):
@@
As the initial hypothesis space $\Theta_{0}\subset\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplanes yields a series of convex polytopes $\Theta_{i}$s, $i=0,1,2,$....
@@ IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
The following assertion states a relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$.
@@
We briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as in Fig. 5 and Fig. 5, the removal of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$. Since the direction of $\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal of the hypothesis space can be relatively small, as in Fig. 5, or large, as in Fig. 5, depending on the specific direction of the correction.
As the specific direction of user correction is hard to assume, a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$. As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathcal{C}_{i}$, as in Fig. 5. We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and volume reduction rate of the hypothesis space in the next section.
Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$. (a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removed volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$. (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
V Implementation Details and Main Results
@@ V-A The Choice of $\boldsymbol{\theta}_{i}$
At each learning iteration $i$, based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large reduction of the hypothesis space for all possible human directional corrections. Thus, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$. One option is to choose the mass center of $\Theta_{i-1}$. However, computing the centroid of a polytope can be #P-Hard [rademacher2007approximating]. For computational efficiency, we choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in $\Theta_{i-1}$, and will later compare its performance with other choices of $\theta_{i}$ in our experiments.
Definition 1 (Maximum Volume inscribed Ellipsoid (MVE)[boyd2004convex]).
@@
$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
@@
where the indicator $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$. Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [boyd2004convex]:
@@
Recall that in Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by hyperplanes (17), and that the initial hypothesis space (11) is a box. By induction, it leads to that any updated hypothesis space $\Theta_{i}$ is a convex polytope, $i=0,1,2,$.... Thus, for any $\Theta_{i-1}$, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$. Then, the choice of $\boldsymbol{\theta}_{i}$ is
@@
The convex optimization (20) can be solved efficiently [cvx].
@@ V-B Result 1: Finite Iteration Convergence
In the Safe MPC Alignment, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
@@
The lemma above asserts an upper bound of the volume of the hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume. With this, we can now establish the finite iteration convergence of the Safe MPC Alignment, stated below.
@@
The above result asserts the maximum learning iteration to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$. In other words, within $K$ rounds of human correction, the robot can successfully learn a safety constraint. The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$. Since the size of $\bar{\Theta}_{H}$ is unknown in reality, $\rho_{H}>0$ can be interpreted as an estimated size of the unknown user intent set $\bar{\Theta}_{H}$. It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated $\boldsymbol{\theta}_{H}$ at convergence.
The finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place for the Safe MPC Alignment. Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a verification for the hypothesis misspecification, where such condition is not met. We will provide such result in the next subsection.
V-C Result 2: Hypothesis Misspecification Certification
Consider hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$. In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4). The hypothesis misspecification can lead to the undesired behavior of the robot [bobu2020quantifying, ghosal2023effect]. The following result states the Safe MPC Alignment can automatically certify the misspecification of $\Theta_{0}$.
@@ V-C Result 2: Hypothesis Misspecification Certification
In the Safe MPC Alignment, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,$.... If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
@@
The above result states when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$. The above theorem inspires us to empirically identify the misspecification by monitoring the position of the MVE center of $\Theta_{i}$ during the learning update. As shown in Fig. 6, if after the maximum iteration K1K-1 in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
Figure 6: When there is misspecification, the Safe MPC Alignment will converge to some point on the boundary of $\Theta_{0}$. So we can use the distance of $\boldsymbol{\theta}_{K-1}$ to the boundary as an indicator for misspecification.
Specifically, if $\Theta_{0}$ is a box as defined in (10) with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$. If $\boldsymbol{\theta}_{K-1}$ satisfies
@@ V-D Detailed Algorithm of Safe MPC Alignment
With the main results on convergence and misspecification, we present the detailed algorithm in Algorithm 1. With $\Theta_{0}$ and a termination threshold $\rho_{H}$ pre-set, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24). Note that during the execution of the MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3. After K1K-1 iterations, the algorithm checks the misspecification by (25).
Initialization: Initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$, $i=1$
Compute maximum iteration count $K$ from $\rho_{H}$ in (24);
@@
     Human restarts robot from safe region;
@@
VI Numerical Examples
In this section, we will evaluate the Safe MPC Alignment using two numerical examples: A pendulum task, where we will illustrate all theoretical aspects of the proposed method, and a thrust-controlled quadrotor navigation task, where we will demonstrate the capability of the method to learn safety constraints in complex 3D environments. Simulated human correction knowing ground-truth safety constraints will be used. We will compare the proposed method with baselines, and ablate the hyperparameters of the algorithm. All numerical experiments are implemented in Python, running on a desktop with Intel i9-13900K CPU. All our experiment codes are open-sourced at https://github.com/asu-iris/Safe-MPC-Alignment.
(a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
(b) After 3 corrections (with $\boldsymbol{\theta}_{4}$)
(c) After 6 corrections (with $\boldsymbol{\theta}_{7}$)
(d) After 14 corrections (with $\boldsymbol{\theta}_{15}$)
Figure 7: The Safe MPC rollout trajectory after different learning interaction. The learned constraint with $\boldsymbol{\theta}_{1}$ is omitted since it covers the whole first quadrant. As learning proceeds, the constraint boundary (orange line) converges to the ground-truth (red line). When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from correction.
@@ VI-A1 Setup
with $\alpha$ being the angle between the pendulum and gravity, $\mathrm{g}$ the gravity constant, $u$ input torque, $m$, $l$ and $d$ the length, inertia, and damping ratio. The system state is $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$, (26) discretized by the Euler method with $\triangle t=0.02\mathrm{s}$. A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each MPC rollout to simulate the state noise. The task is to control the pendulum to a target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$. We pose an unknown global state constraint on pendulum motion:
@@
respectively, with $R=0.1$ and $Q=\mathrm{diag}(25,10)$. To ensure the safe MPC to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state. Thus, we define
@@
Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$. The simulated (synthetic) correction $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first step is taken as in (6)). The correction is probabilistic to simulate the randomness of human providing corrections. It is easy to verify that the simulated $\mathbf{a}$ is valid in term of satisfying the hypotheses in section IV-A.